OpenStudy (fixxer):

Given 3 surfaces, how can you obtain the plot and bounds for the solid made by intersection by all 3 surfaces? S1 is the xy plane, z=0 S2 x^2+y^2=1 S3 z=1-x^2

OpenStudy (tkhunny):

Typically, examine cross sections and understand intersections. x^2 + y^2 = 1 is a cylinder, parallel to the z-axis and centered at (0,0) in the x-y plane. It has a radius of 1. Important points are (1,0,0);(0,1,0);(-1,0,0);(0,-1,0). Neither x, not y can exceed 1, positive or negative. z = 1 - x^2 is parabolic in the x-z plane and runs parallel to the y-axis. A really long tent. This is necessarily less than 1 in the z-direction. Important points are (0,0,1);(1,0,0);(-1,0,0) Give it some thought.

OpenStudy (fixxer):

Can I use cylindrical cordinates? knowing x^2+y^2=1 can be replaced with x=r*cos(t) and y=r*sin(t) also we got z=1-x^2 which can be written \[z=1-r^2\cos^2(t)\] 0<z<1 and t from 0-2Pi and r from 0-1?

OpenStudy (fixxer):

\[\int\limits \!r\,{\rm d}[z={0\ldots 1-{r}^{2} \left( \cos \left( \theta \right) \right) ^{2}},r={0\ldots 1},\theta={0\ldots 2\,\pi}]\] Still cant plot it thought

OpenStudy (tkhunny):

Please do.

OpenStudy (holsteremission):

For the integral, the upper bound on \(z\) needs to be \(1-\cos^2\theta=\sin^2\theta\). No \(r\) needed here, that only applies in the general coordinate conversion.

OpenStudy (fixxer):

Thanks Holster, I need to go thru trig identities again it seems like. My best attempt to plot this thing is hillarious :P

OpenStudy (holsteremission):

Looks good, though that's just the lateral surface of the solid (the part of the cylinder below the parabolic cylinder and above the plane).

OpenStudy (fixxer):

Finally :)

OpenStudy (tkhunny):

Look at you!! Good work. Excellent practice. One word of warning. 3D is as far as your plotting and imagery can go. One day, maybe, you will need a 4D integral. No visualization, there. It all becomes much more abstract. It is good to think about the abstract before you must.

OpenStudy (fixxer):

Thanks. It took quite a bit of time time though xD 4D integral, didn't even know it was a thing xD what on earth are they used for?

OpenStudy (tkhunny):

Oh, there is FAR MORE mathematics out there than we can use to answer the "used for" question.

OpenStudy (holsteremission):

One use could be finding a probability with respect to a multivariate probability distribution that depends on four random variables. Another perhaps more nebulous application might be finding the "volume" (or more generally, measure) of a four dimensional object.